# Lost in Cultural Translation: Do LLMs Struggle with Math Across Cultural Contexts?

### v1

@@ Abstract
Large Language Models (LLMs) have significantly advanced various fields, particularly coding, mathematical reasoning, and logical problem solving. However, a critical question remains: Do these mathematical reasoning abilities persist when LLMs are presented with culturally adapted math problems? Specifically, how do LLMs perform when faced with math problems embedded in cultural contexts that have no significant representation in main stream web-scale AI training data? To explore this, we generated six synthetic cultural datasets from GSM8K, a widely used benchmark for assessing LLMs’ mathematical reasoning skills. While preserving the mathematical logic and numerical values of the original GSM8K test set, we modify cultural elements such as personal names, food items, place names, etc. These culturally adapted datasets provide a more reliable framework for evaluating LLMs’ mathematical reasoning under shifting cultural contexts. Our findings reveal that LLMs struggle with math problems when cultural references change, even though the underlying mathematical structure remains constant. Smaller models exhibit greater performance drops compared to larger models. Interestingly, our results also suggest that cultural familiarity can enhance mathematical reasoning. Even models with no explicit mathematical training but exposure to relevant cultural contexts sometimes outperform larger, mathematically proficient models on culturally embedded math problems. This study highlights the impact of cultural context on the mathematical reasoning abilities of LLMs, underscoring the need for more diverse and representative training data to improve robustness in real-world applications. The benchmark data sets and script for reproducing the results are available at https://github.com/akarim23131/Lost_in_Cultural_Translation
@@ 1 Introduction
Large Language Models (LLMs) have transformed AI, taking on tasks that once seemed to require human intuition. These models have shown remarkable progress in the field of natural language processing, problem solving, question answering, and computer vision tasks (Li et al. [1], Chu et al. [2], Gunter et al. [3], Team et al. [4]). Their ability to interpret and code complex mathematical reasoning has gained a tremendous amount of attention from the research community (Mirzadeh et al. [5]).
However, beneath their impressive capabilities lies a crucial question: Do LLMs truly understand the world’s cultural diversity, or reflect simply the cultural limitation embedded in their training data? More specifically, do they retain their mathematical reasoning when presented with culturally adapted math word problems? How does cultural diversity—or lack thereof—in their training data impact their mathematical reasoning abilities? Studies have consistently highlighted cultural, gender, and sociopolitical biases in LLMs, particularly their Western-centric learning, which affects fairness and adaptability in different linguistic and cultural contexts (Ramesh et al. [6], Ramezani and Xu [7]). Previous research evaluating moral reasoning for LLM in cultures (Ramezani and Xu [7]) further reinforces this concern, showing that these models are predominantly shaped by Western perspectives and struggle in non-Western settings. Research suggests that LLMs do not engage in formal reasoning (Valmeekam et al. [8]) but instead rely on probabilistic pattern-matching to generate outputs based on training data rather than a true conceptual understanding of symbols or ideas (Boix-Adsera et al. [9]). This same mechanism contributes to Western biases in multilingual models, where perspectives from economically dominant English-speaking countries are prioritized over local cultural contexts (Naous et al. [10]). For instance, Arabic text generation in models like ChatGPT and BLOOM reflects Western-influenced narratives more strongly than monolingual models such as AraGPT-2, highlighting how imbalanced training data shapes both reasoning pathways and cultural representations. As a result, LLMs not only fail to engage in genuine reasoning but also exhibit skewed decision-making patterns that reinforce existing biases in global discourse.
Although LLMs generate coherent responses, their logical reasoning remains inconsistent due to fundamental computational limitations in transformer architectures (Li et al. [11]). Structured memory, such as scratchpads, has been proposed to mitigate these limitations, allowing models to handle more complex reasoning tasks (Li et al. [11]). However, this approach requires generating large volumes of tokens to arrive at a solution (Peng et al. [12]), making it computationally expensive. Studies have shown that LLMs struggle with formal logical reasoning, as their outputs are highly sensitive to individual token changes (Jiang et al. [13]). Even minor variations in input can lead to vastly different reasoning paths, resulting in inconsistent problem-solving performance.
@@
The underlying reason for this variability lies in the way transformers process information. Research suggests that a single transformer layer functions similarly to a one-nearest neighbor algorithm (Li et al. [14]), meaning the model’s reasoning process is heavily influenced by the closest matching examples in its training data. Consequently, tasks that require multiple correct token predictions see an exponential decrease in precision as the number of required tokens increases (Shi et al. [15]). The relationship between training frequency and test performance further supports this idea (Razeghi et al. [16]), demonstrating that LLMs rely more on statistical associations than genuine logical inference, limiting their ability to perform consistent reasoning.
@@
LLMs are highly sensitive to input tokens, and even slight changes in their tokenization process can alter their reasoning (Grattafiori et al. [17]). If they have not been exposed to diverse and underrepresented cultural norms and contexts during training, they may tokenize culturally specific prompts differently, potentially leading to shifts in reasoning and varied responses. Furthermore, as the number of tokens increases or tasks become more complex, the probability of accurately predicting tokens decreases (Shi et al. [18]). This suggests that LLMs can easily be distracted by increased complexity. But what happens when they are prompted by culturally adapted math problems, how does this impact their mathematical reasoning? These findings raise fundamental questions about their ability to generalize mathematical reasoning in diverse cultural backgrounds.
@@
To address these questions, we use the GSM8K dataset (Cobbe et al. [19]), a widely recognized benchmark to evaluate the mathematical reasoning capabilities of LLM (Mirzadeh et al. [5]). However, GSM8K has two key limitations. First, given its extensive use, there is a high probability that LLMs have encountered this dataset during training, making it less reliable for assessing their reasoning abilities. Second, it consists of a single set of math problems, lacking any cultural diversity. To overcome these limitations, we modify the GSM8K test set by introducing culturally adapted versions of each question. Specifically, we synthesize six culturally diverse variants from the original GSM8K test set , one for each continent, while preserving the original mathematical logic and numerical values. The only changes involve replacing culturally specific entities such as names, food items, and contextual references with those relevant to each target culture. A detailed explanation of this synthetic dataset generation is provided in Section 3.
@@
We then evaluate 14 LLMs, varying in size and release period, in these culturally adapted versions of GSM8K. The remainder of the paper is structured as follows. Section 2 details the dataset creation process, Section 3 discusses our culture (country) selection process, Section 4 shows detailed performance analysis of LLMs, and Sections 5 present evaluation methods and results.
@@ 2 Dataset Creation
Our dataset creation process and structure are strongly inspired by the GSM symbol dataset introduced by (Mirzadeh et al. [5]). They used the GSM8K test dataset to generate symbolic variants of questions by replacing names with placeholders. This approach enables variations in names while preserving the core mathematical logic, reasoning, and numerical structure of GSM8K. Additionally in their research, numerical values are also adjusted to introduce further complexity to the questions in the GSM8K test set. Similarly, we used the GSM8K test dataset, converting a total of 1,319 questions into symbolic duplicates. However, our approach extends beyond just name substitutions. We systematically identify various cultural entities present in the questions, including person names, food items, clothing, city names, school subjects, common sports, etc. Each identified entity is carefully examined and replaced, ensuring that the questions reflect different cultural contexts without altering their mathematical logic. Importantly, we maintain all the numerical values and the original structure of the questions, making targeted adjustments only to reflect cultural variations. The general data set creation process is shown in Figure 1.
@@ 2.1 Cultural Entities Recognition
Initially, we select a representative sample of 200 questions from the 1,319 questions in the GSM8K dataset and manually identify cultural entities through a detailed human evaluation. Subsequently, we manually create symbolic versions of seven randomly chosen questions from this subset, replacing the identified cultural entities with accurate placeholders. These manually prepared symbolic questions form the basis of a 7-shot prompt constructed for GPT-4o provided in Figure A1 (Appendix A 1).
@@
We repeat this process multiple times to have the final correct output as explained in Appendix 2. Finally, we obtained a clean and standardized symbolic version of each question, with only identified cultural entities and placeholders at their specific locations, as well as a clean set of 55 cultural entities. An example question can be seen in Figure 2(a), and the full list of cultural entities is included in Table A1 (Appendix A 3).
@@ 2.2 Dictionary Creation
After obtaining a clean symbolic version with 55 unique cultural entities, we create a dictionary for each culture. The process of selecting different cultures (countries) is explained in Section 3. This dictionary is primarily constructed using manual web searches, Wikipedia, country-specific websites, and blogs to gather information about each cultural entity. For instance, if a cultural entity is ‘Person name’, we search the Web, Wikipedia, and various country-specific sources to collect the most common names associated with that culture. Similarly, for the entity ‘food items’, we look for traditional foods specific to that culture and so on for the rest of the entities.
@@
The scraping process to fill the placeholder with various cultural entities is performed manually, allowing us to inspect and validate each entry to ensure accuracy. This also gives us the opportunity to review our list of identified cultural entities again, so we can remove any discrepancies. Through this process, we build and populate our dictionary for each culture. The values in the dictionary are then used to replace the placeholders in the symbolic version of our dataset. A snapshot of a dictionary is given in Figure A3 (Appendix A 4).
@@ 2.3 Mapping Rules
One of the final challenges in our dataset creation is ensuring that the mathematical logic of the cultural dataset questions remains consistent with the original GSM8K test dataset, even after replacing placeholders with cultural entities from our dictionary. The objective is to precisely substitute only cultural entities in their designated placeholders while maintaining consistency for repeated entities in their respective positions. This means:
  1. 1.
    If an entity appears multiple times in a question (e.g., the same person name or food item), it should be replaced with the same entity at all occurrences.
  2. 2.
    All of these aspects must be strictly controlled to maintain the logical consistency of the questions.
To achieve this, we create an indexing or mapping rule for each question because every question has a unique structure and different indexing patterns.
@@
Consider the original GSM8K question in Figure 2(a), the identified cultural entities are Lisa, Peter, and chocolate bars. The symbolic question replaces them with placeholders {Person name} and {food items} in the appropriate places. However, the problem arises when replacing placeholders with actual values from the cultural dictionary. We only have one key for person names: {Person name} (which contains multiple names (values) in the dictionary. If placeholders are replaced randomly, there is a risk that the first two placeholders {Person name} get different names, and when Lisa or Peter appears again, they might be replaced with a different name instead of the previously used one. Similarly, {food items} appears four times in the symbolic question, but in the original question, it always refers to “chocolate bars.” If the logic selects different food items from the dictionary at each occurrence, the mathematical logic of the question breaks down.
To fix this issue, we create a mapping or indexing rule for each question, as shown in Figure 2(b).
  1. 1.
    Word Level Tokenization & Placeholder Indexing:
    • First, we tokenize the symbolic version of the question at the word level.
    • We treat each placeholder (e.g., {Person name}) as a single word token rather than separate words.
    • We then identify the exact indices where the placeholders appear.
  2. 2.
    Mapping Back to Original Question:
    • We tokenize the original question at the word level as well and locate the positions where the actual entities appear.
    • We compare these positions with the placeholder indices in the symbolic version.
    • This gives us a one-to-one mapping between placeholders and their original entities.
  3. 3.
    Maintaining Logical Consistency in Replacement:
    • This type of mapping ensures that when {Person name} appears at index 0, it is replaced with Lisa.
    • When {Person name} appears again at index 10, it must be replaced with the same entity (Lisa), not another random name.
    • The same rule applies to {food items}, ensuring that all four instances are replaced consistently with “chocolate bars,” maintaining the original question’s meaning.
This method removes randomness in entity replacement and ensures that the logical structure of the question remains unchanged. It should be noted that although the provided example illustrates the mapping rules using {Person name} and {food items}, these rules apply universally to all entity types present in our dictionary.
@@ 2.4 Replacement
3 Countries Selection
For the initial selection of a country, we chose Pakistan, as the authors have first-hand cultural knowledge and lived experience in this region. This allowed us to ensure a more accurate and contextually relevant adaptation of the dataset. Since a deep understanding of cultural nuances is essential for meaningful modifications, selecting a country with which the authors are familiar provides a reliable foundation for this study. Beyond this, we established a structured selection criterion for additional countries. We aimed for broad representation by choosing one country from each continent (excluding Antarctica due to the absence of a permanent population) and prioritizing underrepresented and economically disadvantaged countries. These countries were chosen to reflect diverse cultural and socioeconomic contexts that may lack strong representation in AI and technology. By combining personal expertise with a systematic selection process, we strive to create a more comprehensive and meaningful evaluation of cultural biases in LLMs. We used four indicators given below to define underrepresented and economically disadvantaged countries.

### v2

@@ Abstract
We demonstrate that large language models’ (LLMs) mathematical reasoning is culturally sensitive: testing 14 models from Anthropic, OpenAI, Google, Meta, DeepSeek, Mistral, and Microsoft across six culturally adapted variants of the GSM8K benchmark, we find accuracy drops ranging from 0.3% (Claude 3.5 Sonnet) to 5.9% (LLaMA 3.1-8B) when math problems are embedded in unfamiliar cultural contexts, even when the underlying mathematical logic remains unchanged. These statistically significant performance reductions (p << 0.01, confirmed through McNemar tests) reveal that mathematical reasoning in LLMs is not culturally neutral.
To create these variants for Haiti, Moldova, Pakistan, Solomon Islands, Somalia, and Suriname, we systematically replaced cultural entities (names, foods, places, etc.) in 1,198 GSM8K questions while preserving all mathematical operations and numerical values. Validation through 130 randomly sampled questions with two independent human annotators (73.8% and 79.2% accuracy) confirmed entity recognition quality, with discrepancies manually corrected. Our quantitative error analysis of 18,887 instances reveals that while explicitly cultural errors (currency: 4.0%, relationships: 4.5%) represent a minority, cultural adaptation affects broader reasoning patterns, with mathematical reasoning errors comprising 54.7% and calculation errors 34.5% of failures.
Interestingly, cultural familiarity can enhance performance: Mistral Saba, despite lacking explicit mathematical training, outperforms some larger models on Pakistan-adapted problems due to Middle Eastern and South Asian training data exposure. Our analysis suggests multiple mechanisms contribute to performance degradation, including tokenization differences, unfamiliar cultural references disrupting reasoning pathways, and training data biases toward Western contexts. This study underscores the need for more diverse training data to ensure robust LLM performance across global contexts.
@@ 1 Introduction
Large Language Models (LLMs) have transformed AI, taking on tasks that once seemed to require human intuition. These models have shown remarkable progress in the field of natural language processing, problem solving, question answering, and computer vision tasks (Li et al.,, 2025), (Chu et al.,, 2024), (Gunter et al.,, 2024), (Team et al.,, 2023). Their ability to interpret and code complex mathematical reasoning has gained a tremendous amount of attention from the research community (Mirzadeh et al.,, 2024).
However, beneath their impressive capabilities lies a crucial question: Do LLMs truly understand the world’s cultural diversity, or reflect simply the cultural limitation embedded in their training data? More specifically, do they retain their mathematical reasoning when presented with culturally adapted math word problems? How does cultural diversity—or lack thereof—in their training data impact their mathematical reasoning abilities? Studies have consistently highlighted cultural, gender, and sociopolitical biases in LLMs, particularly their Western-centric learning, which affects fairness and adaptability in different linguistic and cultural contexts (Ramesh et al.,, 2023; Ramezani and Xu,, 2023). Hovy and Spruit, (2016) argue that such biases arise because NLP systems reflect the social and cultural characteristics of their training data, often disadvantaging underrepresented populations. Prior work has further shown that surface-level changes in representation, such as names or character roles, can systematically affect model outputs even when the underlying task structure remains constant (Lucy and Bamman,, 2021), and that models exhibit culturally skewed tendencies in multilingual settings (Naous et al.,, 2024).
@@
Beyond cultural bias, LLMs exhibit fundamental limitations in formal reasoning. Research suggests that LLMs do not engage in formal reasoning (Valmeekam et al.,, 2023) but instead rely on probabilistic pattern-matching to generate outputs based on training data rather than a true conceptual understanding of symbols or ideas (Boix-Adsera et al.,, 2024). Studies have shown that LLMs struggle with formal logical reasoning, as their outputs are highly sensitive to individual token changes (Jiang et al.,, 2024). A single transformer layer functions similarly to a one-nearest neighbor algorithm (Li et al.,, 2024), meaning the model’s reasoning process is heavily influenced by the closest matching examples in its training data. The relationship between training frequency and test performance further supports this idea (Razeghi et al.,, 2022), demonstrating that LLMs rely more on statistical associations than genuine logical inference.
@@
These reasoning limitations intersect with cultural context in a specific way. LLMs are highly sensitive to input tokens, and even slight changes in their tokenization process can alter their reasoning (Grattafiori et al.,, 2024). If they have not been exposed to diverse and underrepresented cultural norms and contexts during training, they may tokenize culturally specific prompts differently, potentially leading to shifts in reasoning and varied responses. Furthermore, as the number of tokens increases or tasks become more complex, the probability of accurately predicting tokens decreases (Shi et al.,, 2023).
@@
To address these questions, we use the GSM8K dataset (Cobbe et al.,, 2021), a widely recognized benchmark to evaluate the mathematical reasoning capabilities of LLM (Mirzadeh et al.,, 2024). However, GSM8K has two key limitations. First, given its extensive use, there is a high probability that LLMs have encountered this dataset during training, making it less reliable for assessing their reasoning abilities. Second, it consists of a single set of math problems, lacking any cultural diversity. To overcome these limitations, we modify the GSM8K test set by introducing culturally adapted versions of each question. Specifically, we synthesize six culturally diverse variants from the original GSM8K test set , one for each continent, while preserving the original mathematical logic and numerical values. The only changes involve replacing culturally specific entities such as names, food items, and contextual references with those relevant to each target culture. We then evaluate 14 LLMs, varying in size and release period, in these culturally adapted versions of GSM8K test set.
@@
A detailed comparison of this work with prior related research is provided in Section 2. The remainder of the paper is structured as follows. Section 3 details the dataset creation process, Section 4 discusses our culture (country) selection process, Section 5 shows detailed performance analysis of LLMs, and Sections 6 presents the conclusion.
2 Related Work
A growing body of research has documented that NLP systems reflect the social and cultural characteristics of their training data, often disadvantaging underrepresented populations (Hovy and Spruit,, 2016), with fairness research remaining concentrated predominantly on English and high-resource languages (Ramesh et al.,, 2023; Ramezani and Xu,, 2023). Two lines of work are particularly relevant to our study. Lucy and Bamman, (2021) demonstrate that surface-level changes in representation, specifically character gender in narrative prompts, systematically affect GPT-3 outputs in topics and power dynamics, even when the task structure remains constant. Naous et al., (2024) introduce CAMeL, a dataset contrasting Arab and Western cultural entities, showing that both multilingual and Arabic monolingual LMs exhibit significant bias towards Western entities across tasks including story generation, NER, and sentiment analysis. While these studies establish that cultural context shapes model outputs, both evaluate open-ended tasks without objectively verifiable correct answers, and neither examines mathematical reasoning or spans multiple underrepresented cultures across continents. Our work addresses these gaps by evaluating 14 LLMs on culturally adapted math problems with ground truth answers across six underrepresented cultural contexts spanning six continents.
A related line of research examines the sensitivity of LLM reasoning to surface-level input changes. Jiang et al., (2024) introduce a hypothesis-testing framework demonstrating that LLMs rely on token bias rather than genuine reasoning, by systematically altering tokens such as names and quantifiers in logical fallacy problems while preserving the underlying logic, they show that performance drops significantly, confirmed through McNemar testing. Similarly, Shi et al., (2023) introduce GSM-IC, a variant of GSM8K augmented with artificially inserted irrelevant sentences, finding that model performance dramatically decreases when such distractors are present. Both studies establish that LLMs are sensitive to surface-level contextual changes in reasoning tasks. However, the perturbations in both cases are arbitrary or artificial, designed specifically to expose brittleness, and neither study introduces a cultural dimension. Our central finding is that cultural unfamiliarity disrupts mathematical reasoning even when the underlying problem difficulty and numerical logic remain entirely unchanged, suggesting that model performance on mathematical reasoning tasks may be sensitive to the cultural framing of problems, rather than purely to their mathematical structure.
The GSM8K benchmark introduced by Cobbe et al., (2021) has become the standard for evaluating mathematical reasoning in LLMs, comprising grade-school level math word problems requiring multi-step arithmetic reasoning. While LLM performance on GSM8K has improved substantially, Mirzadeh et al., (2024) demonstrate through their GSM-Symbolic benchmark that this improvement may not reflect genuine reasoning advances, by substituting names and numerical values using symbolic templates, they show that model performance exhibits noticeable variance across different instantiations of the same problem. Our work is directly inspired by this approach of controlled substitution while preserving mathematical logic. However, where GSM-Symbolic varies mathematical surface features within a single Western English cultural context, we vary cultural entities across six real-world cultural contexts spanning six continents. Importantly, our findings reveal that this relationship is bidirectional: while unfamiliar cultural contexts degrade mathematical reasoning performance, cultural familiarity can enhance it, a pattern observed when models with regional training data exposure outperform larger models on culturally familiar variants. This bidirectional effect has no analogue in prior benchmark work. Additionally, research on tokenization fairness suggests that culturally specific vocabulary may be processed less efficiently by models trained predominantly on English text (Petrov et al.,, 2023), (Dang et al.,, 2024), representing one potential contributing factor to the performance differences we observe.
@@ 3 Dataset Creation
Our dataset creation process and structure are strongly inspired by the GSM symbolic dataset introduced by (Mirzadeh et al.,, 2024). They used the GSM8K test dataset to generate symbolic variants of questions by replacing names with placeholders. This approach enables variations in names while preserving the core mathematical logic, reasoning, and numerical structure of GSM8K. Additionally in their research, numerical values are also adjusted to introduce further complexity to the questions in the GSM8K test set. Similarly, we used the GSM8K test dataset, converting a total of 1,319 questions into symbolic duplicates. However, our approach extends beyond just name substitutions. We systematically identify various cultural entities present in the questions, including person names, food items, clothing, city names, school subjects, common sports, etc. Each identified entity is carefully examined and replaced, ensuring that the questions reflect different cultural contexts without altering their mathematical logic. Importantly, we maintain all the numerical values and the original structure of the questions, making targeted adjustments only to reflect cultural variations. The general data set creation process is shown in Figure 1.
@@ 3.1 Cultural Entities Recognition
Initially, we select a representative sample of 200 questions from the 1,319 questions in the GSM8K dataset and manually identify cultural entities through a detailed human evaluation. Subsequently, we manually create symbolic versions of seven randomly chosen questions from this subset, replacing the identified cultural entities with accurate placeholders. These manually prepared symbolic questions form the basis of a 7-shot prompt constructed for GPT-4o provided in Figure A1 (Appendix A).
@@
To validate the quality of GPT-4o’s cultural entity recognition, we conduct a systematic evaluation on 130 randomly sampled questions, representing approximately 11% of the culturally adaptable dataset. Two independent human annotators manually review each question to verify whether GPT-4o correctly identified all cultural entities without omissions or errors. The annotators achieve question-level accuracy scores of 73.8% and 79.2%, respectively, indicating that GPT-4o successfully identified all entities in approximately three out of four questions. We adopted question-level accuracy as our validation metric, as our dataset required every question to have all cultural entities correctly identified, a single missed cultural entity breaks the logical consistency of the entire question, making it unusable regardless of how many other entities were correctly identified. All identified discrepancies from the validation sample were manually reviewed and corrected. Importantly, the correction process was not limited to the sampled questions, informed by the error patterns identified during human validation, we conducted multiple iterative correction cycles across the full dataset, combining GPT-4o automated evaluation with human inspection drawing on our cultural familiarity with the target regions. A detailed description of the validation methodology and results is provided in Appendix A Section 2 as well as the broader multi-layer correction process applied to the full dataset. The iterative correction process for placeholder naming inconsistencies is described in Appendix A Section 3.
Finally, we obtain a clean and standardized symbolic version of each question, with only identified cultural entities and placeholders at their specific locations, as well as a clean set of 54 cultural entity types. An example question can be seen in Figure 2(a), and the full list of cultural entities is included in Table A2 (Appendix A).
@@ 3.2 Dictionary Creation
After obtaining a clean symbolic version with 54 cultural entity types, we create a dictionary for each culture. The process of selecting different cultures (countries) is explained in Section  4. This dictionary is primarily constructed using manual web searches, Wikipedia, country-specific websites, and blogs to gather information about each cultural entity. For instance, if a cultural entity is ‘Person name’, we search the Web, Wikipedia, and various country-specific sources to collect the most common names associated with that culture. Similarly, for the entity ‘food items’, we look for traditional foods specific to that culture and so on for the rest of the entities.
@@
The scraping process to fill the placeholder with various cultural entities is performed manually, allowing us to inspect and validate each entry to ensure accuracy. This also gives us the opportunity to review our list of identified cultural entities again, so we can remove any discrepancies. Through this process, we build and populate our dictionary for each culture. The values in the dictionary are then used to replace the placeholders in the symbolic version of our dataset. A snapshot of a dictionary is given in Figure A3 (Appendix A ).
@@ 3.3 Mapping Rules
One of the final challenges in our dataset creation is ensuring that the mathematical logic of the cultural dataset questions remains consistent with the original GSM8K test dataset, even after replacing placeholders with cultural entities from our dictionary. The objective is to precisely substitute only cultural entities in their designated placeholders while maintaining consistency for entities that appear multiple times within the same question; for example, if the same person name appears multiple times, it must be replaced with the same culturally adapted name at every occurrence. To achieve this, we create an indexing or mapping rule for each question, as illustrated in Figure 2(b), which establishes a one-to-one mapping between each placeholder instance and its corresponding original entity. This method removes randomness in entity replacement and ensures that the logical structure of every question remains unchanged across all entity types in our dictionary.
@@ 3.4 Replacement
4 Country Selection
For the initial selection of a country, we chose Pakistan, as the authors have first-hand cultural knowledge and lived experience in this region. This allowed us to ensure a more accurate and contextually relevant adaptation of the dataset. Since a deep understanding of cultural nuances is essential for meaningful modifications, selecting a country with which the authors are familiar provides a reliable foundation for this study. Beyond this, we established a structured selection criterion for additional countries. We aimed for broad representation by choosing one country from each continent (excluding Antarctica due to the absence of a permanent population) and prioritizing underrepresented and economically disadvantaged countries. These countries were chosen to reflect diverse cultural and socioeconomic contexts that may lack strong representation in AI and technology. By combining personal expertise with a systematic selection process, we strive to create a more comprehensive and meaningful evaluation of cultural biases in LLMs. We used three indicators given below to define underrepresented and economically disadvantaged countries.